Grade 12th passGeneral Physics

three identical rods of length l and mass m are connected to form U shape . Find distance of centre of mass from the centre o .

Chanchal
10 Years agoGrade 12th pass

1 Answer

Vikas TU
10 Years ago
Center of mass of the system = 
if the lower postion of leftmot rod is assumed to be origin then,
xcom = (m*0 + m*l/2 + m*l)/3m = l/2
ycom = (m*l/2 + m*0 + m*l/2)/3m = l/3
 
Hence com for the system is at :
(l/2, l/3)
find the distance of this point to centre 0 by usong distance formulae.
 
U have not mentioned where the point O is located?